# Reward-Safety Balance in Offline Safe RL via Diffusion Regularization

### v1

@@ Appendix B Supplemental Materials for Algorithm 1
We outline the Safe Adaptation step in Algorithm 1, specifically detailed in Algorithm 2 and Algorithm 3. Our main theorem (Theorem 3.1) is derived based on the safety adaptation procedure described in Algorithm 3. Our diffusion regularization method is compatible with both CRPO Xu et al. [2021] and the gradient manipulation method Gu et al. [2024a], as implemented in Algorithms 2 and 3. Both algorithms aim to switch dynamically between reward optimization and cost optimization.
@@
The key distinction lies in their approach to handling scenarios where the cost is close to the cost limit threshold. In Algorithm 2, the process focuses on optimizing one of the two objectives based on predefined conditions. In contrast, Algorithm 3 incorporates gradient manipulation in these scenarios, further stabilizing the training process by addressing conflicts between the objectives. Theorem 3.1 specifically considers the diffusion regularization algorithm equipped with the safe adaptation procedure outlined in Algorithm 3. The fundamental difference between the two algorithms lies in the criteria for switching between reward and cost optimization objectives.
Algorithm 2 CRPO Adaptation
1:Dataset $\mathcal{D}^{\mu}$
2:Slack variable $h^{+}$, $h^{-}$ and cost limit $l$
3:Procedure: CRPOAdaptation$(\pi_{\theta},\epsilon_{\psi},Q^{\pi_{\theta}}_{r},Q^{\pi_{\theta}}_{c},h^{+},h^{-})$
4:if $Q^{\pi,\text{UCB}}_{c}(\rho)\leq l+h^{+}$ then
5:     Optimize reward by solving Eq. (20)
6:else
7:     Ensure safety by solving Eq. (21)
8:end if
9:end procedure
@@
1:Dataset $\mathcal{D}^{\mu}$
2:Slack variable $h^{+}$, $h^{-}$ and cost limit $l$
3:Procedure: SafeAdaptation$(\pi_{\theta},\epsilon_{\psi},Q^{\pi_{\theta}}_{r},Q^{\pi_{\theta}}_{c},h^{+},h^{-})$
4:if $Q^{\pi,\text{UCB}}_{c}(\rho)\leq l-h^{-}$ then
5:     Optimize reward by solving Eq. (20)
6:else if $Q^{\pi,\text{UCB}}_{c}(\rho)\leq l+h^{+}$ then
7:     Compute $g_{r}$ and $g_{c}$ with Eq. (23)
8:     Gradient manipulation to obtain $g$ with Eq. (12)
9:else
10:     Ensure safety by solving Eq. (21)
11:end if
12:end procedure
B.1 Hybrid Extension
@@
Appendix C Further Discussion on Safe Reinforcement Learning
@@
$c(s_{t},a_{t})=0,\forall t\geq 0.$ (50)
Safe Constraints. A variety of works focus on the soft constraint setting, the safe constraint restricts the cost limit below a certain limit $l$, either within an episode or being step wise. Which can either can be episodic limit as
@@ C.1 Discussion on Settings
Appendix D Details of Experiments
D.1 Choice of Policy Class
@@
D.2 Discussion on Hyperparameters
Choice of Slack Variable. We define the slack bounds based on the normalized cost, ensuring that reward optimization is performed when $V_{\text{normalized}}\leq 1-h^{-}$ and cost optimization is triggered only when $V_{\text{normalized}}\geq 1+h^{+}$. Over the course of training, we progressively reduce the values of $h^{-}$ and $h^{+}$, eventually shrinking them to zero. For CRPO adaptation, we set $h=0.2$. In our gradient manipulation method, as outlined in Algorithm 3, we initialize both $h^{-}$ and $h^{+}$ to $0.2$ and shrink the slack variables’ value by iterations.
Hyperparameters BC BCQ-Lag BEAR-Lag CDT CPQ COptiDICE Ours
Device Cuda Cuda Cuda Cuda Cuda Cuda Cuda
Batch Size 512 512 512 2048 512 512 256
Update Steps 100000 100000 300000 100000 100000 100000 2050
Eval Steps 2500 2500 2500 2500 2500 2500 1025
Threads 4 4 4 6 4 4 4
Num workers 8 8 8 8 8 8 8
Actor Architecture(MLP) [256,256] [256,256] [256,256] \ [256,256] [256,256] [256,256]
Critic Architecture(MLP) \ [256,256] [256,256] \ [256,256] [256,256] [256,256]
Actor Learning rate .001 .001 .001 .001 .001 .001 .0006
Critic Learning rate \ .001 .001 \ .001 .001 .0006
Episode Length 1000 1000 1000 1000 1000 1000 1000
$\gamma$ 1.00 0.99 .99 .99 .99 .99 .99
$\tau$005 .005 .005 .005 .005 .005 .005
$h^{+}$2
$h^{-}$2
PID(optional for DRCORL) \ [.1,.003, .001] [.1,.003,.001] \ \ \ [.1,.003,.001]
$E$ 4
$k$ 2.0
$\alpha$2
Remark D.1.
D.3 Experiment Environments
As depicted in Figure 5, this is the experiment environment we use, which is mujoco-Gym environments. We also attach the training curve for different tasks in this section. For practical implementation, we may consider three types of policy class $\Pi$:
@@ D.3 Experiment Environments
Figure 5: Mujoco-Gym Environments

### v2

@@ Appendix B Supplemental Materials for Algorithm 1
B.1 Clarification for the Main Body
Inconsistency in critic update. In the training process, we specifically adopt a conservative estimation for the cost critic, the differing update strategies for $Q^{\pi}_{r}$ and $Q^{\pi}_{c}$ are intentional and reflect their distinct roles in our framework. The cost critic $Q^{\pi}_{c}$ plays a critical role in enforcing safety, so we adopt a pessimistic update to penalize uncertain or unsafe state-action pairs—particularly in low-coverage areas—consistent with prior work on conservative critics in offline RL.
In contrast, the reward critic $Q^{\pi}_{r}$ guides policy optimization within the safe region. Applying pessimism here may lead to overly conservative behavior, reducing performance. We therefore use a standard update to maintain a balance between safety and reward. Our gradient manipulation mechanism relies on $Q^{\pi}_{c}$ to modulate the influence of each objective, making a pessimistic cost critic essential for reliable safety adaptation. We appreciate the concern about value overestimation and will explore robust reward estimation techniques in future work.
Pretraining stage. We pretrain the diffusion model to accurately capture the behavioral policy from the offline dataset, enabling it to serve as a stable regularizer for the Gaussian policy during training. Unlike diffusion-based planning methods, our approach uses the diffusion model’s score function to penalize out-of-distribution actions, improving stability.
After pretraining, the diffusion model remains fixed. However, updating the critics during training is crucial, as they guide the gradient manipulation mechanism to balance reward maximization and safety. This separation ensures effective and stable policy optimization throughout training.
Gradient Manipulation Stage. the gradient manipulation method employed in our approach differs fundamentally from the naive weighted average of the two objectives. Specifically, we dynamically adjust gradient weights based on the extent of safety constraint violations, guided by slack variables. When the safety threshold is significantly violated, parameter updates are driven primarily by the gradient of the safety objective. Conversely, when the current policy strictly adheres to safety constraints, updates are performed using only the gradient of the performance objective. In intermediate scenarios, we judiciously combine both gradients.
To address the potential instability , our approach carefully assigns weights according to the angle between gradients, as detailed in Equation (12) of the main body. Instability typically arises when the angle between gradients exceeds $90^{\circ}$. Our weighting strategy effectively mitigates gradient conflicts, preventing gradient degradation and enhancing the stability and reliability of the optimization process. Additionally, our method does not require the reward and cost critics to share parameters. Instead, we temporarily freeze and save the policy parameters to implement gradient manipulation. For example, given two separate critics, $Q^{\pi}_{r}$ and $Q^{\pi}_{c}$, we first store the initial parameters of the policy $\pi_{0}$. We then independently compute updates using each critic. By comparing the updated policy parameters with the initial parameters, we obtain the gradient values required for manipulation. Through this scenario-based approach to parameter updates, we effectively ensure training stability.
B.2 Safety Adaptation Step
@@
We outline the Safe Adaptation step in Algorithm 1, specifically detailed in Algorithm 2. Our main theorem (Theorem 3.1) is derived based on the safety adaptation procedure described in Algorithm 2. Our diffusion regularization method is compatible with both CRPO Xu et al. [2021] and the gradient manipulation method Gu et al. [2024a], as implemented in Algorithm 2. Both algorithms aim to switch dynamically between reward optimization and cost optimization.
@@
The key distinction lies in their approach to handling scenarios where the cost is close to the cost limit threshold. In Algorithm Algorithm 2 incorporates gradient manipulation in these scenarios, further stabilizing the training process by addressing conflicts between the objectives. Theorem 3.1 specifically considers the diffusion regularization algorithm equipped with the safe adaptation procedure outlined in Algorithm 2. The fundamental difference between the two algorithms lies in the criteria for switching between reward and cost optimization objectives.
@@
Dataset $\mathcal{D}^{\mu}$
Slack variable $^{+}, hh^{-}$ and cost limit $l$
[…]
if $Q^{\pi,\text{UCB}}_{c}(\rho)\leq l-h^{-}$ then
  Optimize reward by solving Eq. (19)
else if $Q^{\pi,\text{UCB}}_{c}(\rho)\leq l+h^{+}$ then
  Compute $g_{r}$ and $g_{c}$ with Eq. (21)
  Gradient manipulation to obtain $g$ with Eq. (12)
else
  Ensure safety by solving Eq. (20)
end if
end procedure
B.3 Hybrid Extension
@@
B.4 Further Discussion on Safe Reinforcement Learning
@@
$c(s_{t},a_{t})=0,\forall~t\geq 0.$ (45)
Soft Constraints. A variety of works focus on the soft constraint setting, the safe constraint restricts the cost limit below a certain limit $l$, either within an episode or being step wise. Which can either can be episodic limit as
@@ B.5 Discussion on Settings
Appendix C Details of Experiments
C.1 Further Ablation Study
In the initial training phase, we set $\beta$ to a low value (starting at 0.04) to ensure the policy remains close to the behavioral policy, facilitating a stable foundation and minimizing out-of-distribution actions. As training progresses, we gradually increase $\beta$, eventually reaching 1.0, to place greater emphasis on optimizing the critics. This linear scheduling allows for a smooth transition from imitation to optimization, balancing exploration and exploitation effectively. Improper tuning of $\beta$ can impact performance:
- If $\beta$ increases too rapidly: The policy may deviate prematurely from the behavioral policy, leading to instability and potential safety violations due to insufficient grounding in safe behaviors.
- If $\beta$ increases too slowly: The policy may remain overly conservative, limiting performance improvements and resulting in suboptimal reward outcomes.
Therefore, a carefully planned $\beta$ schedule is crucial to balance safety and performance. Our linear approach has demonstrated effectiveness across various tasks.
Choice of Slack Variable. We set slack bounds relative to the normalized cost so that reward maximization applies when $V_{\text{normalized}}\leq 1-h^{-}$ and cost minimization triggers when $V_{\text{normalized}}\geq 1+h^{+}$. During training, both $h^{-}$ and $h^{+}$ are linearly decayed from 0.2 to zero. An ablation study on the impact of slack values is shown in Figure 4. We tested on different initial values for $h^{+}=h^{-}=h\in\{0.1,0.3,0.5,0.7\}$.
Figure 4: Slack Ablation
Choice of cost limit. We evaluated our algorithm on OfflineCarRun-v0 using three cost limits ($l\in\{10,20,30\}$) with five random seeds each. Figure 5 presents the normalized reward and cost. The normalized reward remains consistent across different limits, and the learned policy reliably keeps the cost below the safety threshold.
Figure 5: Cost Limit Ablation
Hyperparameters BC-Safe BEARL BCQ-Lag CPQ COptiDICE CDT CAPS CCAC Ours
Device Cuda Cuda Cuda Cuda Cuda Cuda Cuda Cuda Cuda
Batch Size 512 512 512 512 512 2048 512 512 256
Update Steps 100000 300000 100000 100000 100000 100000 100000 100000 2050
Eval Steps 2500 2500 2500 2500 2500 2500 2500 2500 1025
Threads 4 4 4 4 4 6 4 4 4
Num workers 8 8 8 8 8 8 8 8 8
Actor Architecture(MLP) [256,256] [256,256] [256,256] [256,256] [256,256] \ [256,256] [256,256] [256,256]
Critic Architecture(MLP) \ [256,256] [256,256] [256,256] [256,256] \ [256,256] [256,256] [256,256]
Actor Learning rate .001 .001 .001 .001 .001 .001 .001 .001 .0006
Critic Learning rate \ .001 .001 .001 .001 \ .001 .001 .0006
Episode Length 1000 1000 1000 1000 1000 1000 1000 1000 1000
$\gamma$ 1.00 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99
$\tau$ .005 .005 .005 .005 .005 .005 .005 .005 .005
$h^{+}$ \ \ \ \ \ \ \ \ .2
$h^{-}$ \ \ \ \ \ \ \ \ .2
PID \ [.1,.003,.001] [.1,.003,.001] \ \ \ \ \ [.1,.003,.001]
$E$ \ \ \ \ \ \ \ \ 4
$k$ \ \ \ \ \ \ \ \ 2.0
$\alpha$ \ \ \ \ \ \ \ \ .2
Remark C.1.
C.2 Choice of Policy Class
@@
C.3 Training Details
As illustrated in Figure 6, our experiments are conducted in Safety–Gym environments and Bullet-Gymnasium environments. In this section, we present the training curves for each task. For every task, we run five independent seeds and plot, at each training step, the mean and standard deviation of both reward and cost.
@@ C.3 Training Details
Figure 6: Safety-Gym Environments
@@
Figure 13: OfflineBallCircle Reward
Figure 14: OfflineBallCircle Cost
Figure 15: OfflineCarCircle Reward
Figure 16: OfflineCarCircle Cost
Figure 17: OfflineCarRun Reward
Figure 18: OfflineCarRun Cost
Figure 19: OfflineCarGoal1 Reward
Figure 20: OfflineCarGoal1 Cost
Figure 21: OfflineCarGoal2 Reward
Figure 22: OfflineCarGoal2 Cost
Figure 23: OfflinePointGoal1 Reward
Figure 24: OfflinePointGoal1 Cost
Figure 25: OfflinePointGoal2 Reward
Figure 26: OfflinePointGoal2 Cost
Figure 27: OfflinePointCircle1 Reward
Figure 28: OfflinePointCircle1 Cost
Figure 29: OfflinePointButton1 Reward
Figure 30: OfflinePointButton1 Cost